### v1

2 SRON Netherlands Institute for Space Research, Landleven 12, 9747 AD Groningen, Netherlands
3 Kapteyn Astronomical Institute, University of Groningen, The Netherlands
4 Pontificia Universidad Catolica de Chile, San Joaquin, Chile
5 SRON Netherlands Institute for Space Research, Niels Bohrweg 4, 2333 CA Leiden, The Netherlands
6 Faculty of Electrical Engineering, Mathematics and Computer Science, Delft University of Technology, Mekelweg 4, 2628 CD Delft, The Netherlands
7 I.Physikalisches Institut, Universität zu Köln, Zülpicher Straße 77, 50937 Cologne, Germany
8 Center for Astrophysics Harvard & Smithsonian, 60 Garden St, Cambridge, MA 02138, U.S.A.
11email: nireyes@mpifr-bonn.mpg.de
AMKID - a large KID-based camera at the APEX telescope
The thermal emission at sub-millimeter wavelengths carries unique information in many astronomical applications ranging from disks and planet formation around young stars, to galaxy evolution studies at cosmological distances. Advancing on the mapping speed to detect this faint emission in ground-based astronomy has been a technical challenge for decades.
The APEX Microwave Kinetic Inductance Detector (AMKID) camera was designed to accomplish this task. The instrument is a wide field-of-view camera based on kinetic inductance detectors. It is installed on the 12 meter APEX telescope in Chile at 5.100 meters above see level. The instrument operates dual color, covering simultaneously the 350 GHz and 850 GHz atmospheric windows. It has a large field-of-view of 15.3’x15.3’, and an unprecedented number of pixels: 13.952 detectors in the high frequency band and 3.520 detectors in the low frequency band. Here we present a complete description of the instrument design and construction together with results of the successful low frequency array (LFA) commissioning campaign executed during the last year.
The LFA performance is in good agreement with design parameters, with detector sensitivity of 2.2 mK and diffraction limited beam sizes of 17.0”. On-sky measurements demonstrate a sensitivity of 70-90 mJy per detector when operating under good atmospheric conditions (PWV below 1.0mm). With this performance the LFA regularly achieve a mapping sensitivity of 25 mJy when mapping a square degree in an hour.
AMKID on APEX with its dual color observing capabilities, high sensitivity, large field-of-view and high angular resolution holds the promise to open a new range of science with the APEX telescope.
Key Words.:
Incoherent detectors – astronomical instruments – experimental methods1 Introduction
Until its de-commissioning in 2020 the wide-field direct detection camera LABOCA (Kreysa_2003; Siringo2009), operating up to 295 horn-coupled thermistors in the 870µm atmospheric window, has been the most successful instrument at the APEX telescope, described in (Guesten_2006), leaving a fantastic legacy with close to 500 scientific publications. LABOCA’s noise-weighted point-source sensitivity per channel of 55 mJy, after removal of correlated sky noise, served as gold standard for subsequent developments, but the underlying technology also limited the number of detectors in the camera’s field of view.
A path to more efficient large-scale observations, towards focal-plane arrays with several kilo-pixels at affordable costs, did open in the early 2000s with the introduction of the kinetic inductance detector (KID); (Day_2003). Since then, a number of KID-based cameras have been deployed for operation in the mm regime, such as NIKA and NIKA2 at the IRAM-30m telescope (Monfardini_2010; NIKA2), or MUSIC at the Caltech Submillimeter Observatory (Maloney_2010). Other instruments, also covering the submillimeter regime, are currently at different stages of development as the TolTec (Wilson_2020), BLAST-TNG (Lourie_2018) or P-CAM (Vaskuri_2025).
Building on these promises MPIfR with SRON have launched an ambitious project to develop and operate a wide-field, KID-based submm-camera for the APEX telescope: the APEX microwave KID camera (AMKID), initially described in Esteras_2015. It is a dual-color instrument, operating in the 870 and 350 m atmospheric window simultaneously, thereby taking advantage of the superb atmospheric conditions at Llano de Chajnantor and the excellent sub-millimeter performance of the telescope. The detectors are lens coupled MKID devices, filling the 15 arcmin field-of-view of the telescope in technically densest possible arrangement, meaning several thousands of pixels detectors per array. Taking advantage of the site, operation of the detectors is close to be background limited.
In this article we report on the overall instrument design and focus on the successful commissioning of the low-frequency camera. The high-frequency complement will follow shortly, as will the opportunity for polarization observations.
For a more detailed discussion on the scientific motivations we refer the reader to the extensive literature on this subjects, which range from disks and planet formation around young stars (e.g. Andrews_2018; Holland_2017), to large scale Milky Way studies (e.g. Schuller_2009; Molinari_2010), or galaxy evolution at cosmological distances (e.g. Weiss_2009; Oliver_2012) just to mention some relevant examples. The promises of mapping large scale structures on southern skies, be it galactic or cosmological, with the powerful state-of-the-art AMKID arrays installed at the APEX on one of the world’s outstanding submm-sites are unique and exciting.
2 Instrument Overview
The AMKID instrument is a wide-field sub-millimeter camera. It operates at 870 m (345 GHz) and 350 m (850 GHz) using its Low Frequency Array (LFA) and High Frequency Array (HFA) respectively. Both arrays are diplexed in polarization by a free-standing wire polarizer and can be used simultaneously. The instrument is designed for operations at the 12 meter APEX telescope, making use of the superb observation conditions at the Chajnantor plateu. Precipitable water vapour (PWV) below 0.5 mm is observed over several weeks a year, allowing regular operation in the 350 m band. In addition the telescope surface quality is regularly monitored and adjusted via holography, keeping it in the range of 15-20 m RMS over the complete observation season.
The instrument optics is designed to make use of the complete field-of-view avalaible at the telescope (15.3’ x 15.3’). The field is sampled with an unprecedented number of pixels: 3,520 pixels in the low frequency band and 13,952 pixels in the high frequency band, using a pixel spacing of 1.2/D = 18” at the LFA and 1.5/D = 9” at the HFA. The instrument optics is near the diffraction limit, with nominal half maximum beam sizes of 17.0” at the LFA and 7.5” at the HFA.
3 Detector technology
KIDs were pioneered in the early 2000s by the group of J. Zmuidzinas at Caltech (see Day_2003). They are superconducting microwave resonators designed to efficiently absorb radiation exceeding their gap frequency, which is 90 GHz for Aluminum. Upon radiation absorption Cooper pairs are broken into single particle excitations called quasiparticles, resulting in an increase in both the (kinetic) inductance and the extremely small resistance of the superconducting film at microwave frequencies. As a result, the resonance frequency decreases which can be read-out very accurately using a single microwave probe tone at the unperturbed resonant frequency.
To eliminate thermal excitations and excess noise KIDs need to be operated far below the superconductor critical temperature, K for Aluminum. At these temperatures all charge transport is achieved by Cooper pairs, which have zero resistance, but a finite (kinetic) inductance associated with the electron mass.
The main advantage of this technology is that individual KIDs can be designed to each have a specific resonance frequency, so that each detector in a KID array is associated with a single resonance dip, as shown in Fig. 3. As is clear from this figure many KIDs can be multiplexed in frequency space and read using a single microwave line, a single cryogenic amplifier, and a single room-temperature back-end. This concept
allows to build large arrays, opposed to other previous technologies where each detector
needed its own wiring, limiting the number of pixels.
The AMKID detectors are NbTiN-Aluminum lens-antenna coupled hybrid KIDs as shown in Fig. 2. These detectors, fabricated on high resistivity silicon, are described in Jansen_2013 for the LFA and in Ferrari_2018a for the HFA. They reach photon noise limited performance for the loading powers relevant for AMKID combined with a high coupling efficiency. The Appendix A offers a more detailed overview on the used detector technology.
To allow for small detector sizes we read-out the KIDs in the frequency band of 4 to 8 GHz. Resonances are designed to operate with depths between 3 and 6 dB and Q factor of the order of 2.1E4 under the
target optical load, as seen in Fig. 3. The monolithic silicon lens arrays are fabricated using laser ablation, with a typical fabrication accuracy of 5 m in form factor. Each lens array is coated with a anti-reflection coating made from Parylene-C, as described in Ji2000. The detector efficiency is estimated using numeric simulations to be 84% and 76% at the low and high frequency bands. Between the lens and the planar antenna we use a capacitive grid which absorbs unwanted in-chip stray light (see Yates_2017).
Each array is composed of four 61x61 mm2 silicon chips with the associated lens arrays as shown in Fig. 5, 11 and 14. The lens arrays and chips are assembled using dedicated alignment tooling and markers on the lens array and chips. A permanent bond is achieved using cyanoacrylate glue. The LFA chips have 880 detectors each, and are read using a single microwave line. On the other side the HFA silicon chips have 3488 pixels and are read using 5 coaxial lines, i.e. around 700 detectors per line.
4 Characterization of the detector arrays
Detector chips were routinely characterized in the laboratory using a dedicated test bed, described in Ferrari_2018b. Measurements are used to verify that the produced detector chip met specifications before integration to the instrument. Here we would like to highlight some relevant results.
The achieved frequency error ratio ([) for the LFA was on the order of using contact lythography. Instead the HFA chips, fabricated in a later stage of the project using electron beam lithography for the aluminum patterning, exhibit errors on the order of . It’s remarkable that this value approaches the best results obtained with chip post-processing presented in Shu2018. As we are closely packing our detectors (880 detectors in 4 GHz bandwidth) we have a typical spacing between resonators of , see measured value in Fig 3.(c), still below the fabrication accuracy. This causes the grouping of resonances observed in Fig. 7. Our calibration scheme recovers most of the affected detectors but a loss of around a 10% in the yield is caused by this phenomena.
An important characteristic of direct detectors is the noise power spectral density (PSD). This gives the detector output noise, which is a combination of a white, photon noise spectrum; and a 1/f term due to two-level systems, as discussed in Gao2008. Our detectors are characterized by a 1/f knee frequency of 0.5 Hz. A lower knee frequency is possible by increasing the detector responsivity, but this would result in broader resonance dips and more resonance frequency collisions per readout line. The current configuration was chosen as a good compromise between the two effects.
5 Optical system
The optical system of AMKID is composed of six reflective mirrors as depicted in Fig. 4. The system was designed to make use of the maximum possible field-of-view at the APEX telescope. The initial optical layout, which was later optimized, consists of two Gaussian beam telescopes with a total magnification of 3.6.
The first Gaussian telescope, composed of mirrors M6 and M5, is located inside the cryostat and has a magnification near unity. The output image plane is used as location for the cryostat window. A cold pupil, at around 8 K, is located between these two mirrors and shapes the beam to the sub-reflector size. A cold baffle structure is used as termination for stray light in the system. Electromagnetic simulations (CST studio suite ®) were used to compute the coupling efficiency of detectors to the camera pupil as 45% for LFA and 55% for HFA. The rather high loss is because of the choice of focal plane sampling, trading off mapping speed versus single pixel performance by oversampling as discussed in Griffin_2002.
The second Gaussian telescope, consisting of the mirrors M3 and M4, is located outside the cryostat. It converts the image plane located at window location to the telescope focal surface. Telescope boundaries, given by the size and shape of the available space in the Cassegrain cabin, requires the use of two additional flat mirrors (F1 and F2) to fold the optic path. The use of fully reflective optics circumvents the otherwise critical anti-reflection coating of lenses for multi-frequency operation.
This initial design was optimized to compensate for optical aberrations over the complete field of view. The optimization was done using geometric optics (Zemax OpticStudio®) by define all mirror surfaces, including the flat folding mirrors, as extended polynomial expressions. The coefficients that define mirror shapes were optimized with the goal of achieving a Strehl ratio at 350 larger than 0.8 over the complete field of view, meanwhile constraining the system to be diffraction limited. A detailed overview of the design process and final results is presented in Esteras_2015.
A Monte Carlo tolerance analysis shows that high precision alignment of the mirrors is required with regard to the crystat body: 0.2mm for critical mirrors as F2 and M4, and 0.5mm for M3 and F1 mirrors. Internal cryogenic mirrors are built as a single block and cannot be adjusted. The analysis show that under this limits the Strehl ratio degradation is lower than 0.95 at 850 GHz. The instrument optics is aligned and regularly verified using a metrology arm (FaroArm®) which is installed as part of the instrument in the telescope Cassegrain cabin.
The optical design was verified using physical optics (GRASP from TICRA) . A set of 9 beams across the array was investigated: the four corners of the field of view, the four centers of individual sub-arrays and the center position or system optical axis. For each beam the current distribution over the mirrors was calculated. The cold-stop and window aperture truncation were computed as virtual currents on these apertures. The LFA numeric simulation includes the telescope sub-reflector and the main 12 meter dish. Given numeric complexity, the HFA simulation does not include the antenna, but instead, a geometric optical model is used to propagate the sub reflector illumination to the main aperture of the telescope. Subsequently Fourier optics is used to transform the E-field distribution on the telescope aperture into the beam shape on sky. Results confirm that the optical system is diffraction limited with a beam size of less than 17” for the LFA and less than 8” for the HFA. The ellipticity of beams in both arrays is lower than 1.1 across the complete field of view.
6 Cryogenic system
The KIDs need to be operated at a temperature low enough that the quasiparticle density is dominated by the expected sky loading power. For AMKID, given the aluminum critical temperature of 1.2 K, this leads to a temperature requirement of 260 mK for the LFA and 280 mK for the HFA. This is achieved by using a commercial three-stages sorption cooler. The first stage of the device is based on a 4He sorption pump and reaches 1.1 K. The subsequent two stages, both based on 3He sorption devices, reach 300 mK and 260 mK under operating conditions. The last stage is used to effectively cool the detectors to the required operating temperature. The holding time of the system under operational conditions is more than 10 hours and recycling of the refrigerator system takes about 2 hours. This allows a typical operation time of 20 hours per day.
Fig. 5 shows a photo of the structure holding the sorption cooler and the detector arrays. The four temperatures stages are compactly assembled via a series of hexapod structures and concentric shields. Superconductive niobium-titanium (NbTi) coaxial cables are used to minimize conductive heat losses between stages. The tight structure and the use of absorbers minimize stray light radiation reaching the detectors.
The base temperature required to operate the sub-K cooler is provided by a dedicated two-stages pulse tube operating at 3.3 K. A second pulse tube is used to cool the cold optics, the inner optical enclosure including the baffle structure, and the 24 cryogenic low-noise amplifiers used for the instrument read-out. The final temperature of this stage is 8 K with a thermal load of around 4 W. The first stages of both pulse tubes are connected to the infrared shield, achieving a typical temperature between 40 and 50 K. Fig 5.b presents a diagram of the cryostat with the main parts described here. The instrument is located in the Cassegrain cabin of the telescope, therefore its mechanical orientation changes with telescope elevation. The system is designed to optimally operate at 50 deg elevation, where the sorption cooler and the pulse tubes are near to vertical orientation. However, all cryogenic subsystems can operate within specifications in the elevation range between 30 deg and 80 deg. Of special concern during operations is the recycle process; maximal efficiency, i.e. maximum holding time, is achieved when the recycle is done in the nominal position, 50 deg elevation. In particular, the nominal park position, 11 deg of elevation, has to be avoided during the recycling process.
7 Filter stack
KIDs are sensitive to any radiation above the superconductor band gap that reaches the inductive transmission line. Therefore, a set of filters is required to define the instrument bandpass. The filter set, which also includes the input cryostat window, also play a relevant role in rejecting infrared and optical radiation.
The cryostat window is located at a highly aberrated image of the focal plane and has a size of 15x15 cm2. The window is built from a 5 mm high density polyethylene (HDPE) plate coated on both sides with a 90 m Zitex® layer acting as anti-reflection coating for the 850 GHz frequency range. Given the relative low refraction index of both materials the reflection losses at 350 GHz are of order 5 only.
The power that enters the cryostat via the window aperture is around 11 W. The spectrum corresponds to a 300 K black body which peaks in the infrared. This radiation is mostly reflected by a 3 THz low-pass infrared filter located at the 50 K shield, allowing operation of the inner 4 K and lower stages. This filter is designed to be unaffected by self heating effects due to its reflective design and its implementation as a single layer structure on a thin substrate, see Tucker2006. The remaining infrared power is filtered by a second 5 mm HDPE sheet coated with Zitex installed just after the reflective filter.
A second filtering step, a 1 THz multilayer low-pass filter, is installed at the 8 K optical enclosure . Thermometers are used to monitor the temperatures on the filter edges (130 K for the HDPE and reflective filter, 17 K for the 1 THz low-pass). This information is used to feed our model of the filter stack described in the next subsection. The large difference between the cold-head temperatures and filters near the window is caused by the large distance separating the two, the large infrared load on our cryostat, and the thermal conduction of the radiation shield structure.
After the window and infrared filtering stage the radiation enters the optical cavity, at a nominal temperature of 8 K, where M5, M6, and the cold stop are installed. The cavity contains a cold-baffle structure that terminates warm stray radiation that enters into the cryostat from the instrument cabin environment. It also provides a cold termination for detector spill-over. After the optical cavity radiation reach the detector area where the free-standing wire grid diplexes both arrays in polarization. One linear polarization is reflected towards the LFA meanwhile the other polarization is transmitted to the HFA. On both arrays a set of three filters is used to define the pass-band. The first two, attached to the 4K and 1K stages, are low pass filters. The last filter, located at the 300 mK stage, is a bandpass that define the operation band of the arrays. Figure 1 presents the bandpass of the filter stack, including cryostat window, used in each array. This information is summarized in Table 1.
| Parameter | Central frequency | Bandwidth | Transmission |
| LFA | 350 GHz | 38 GHz | 56.5% |
| HFA | 850 GHz | 123 GHz | 41.4% |
7.1 Filter self-heating model
As part of the design of the instrument a model of the filter stack had been established. For each filter the transmitted, reflected, and emitted power were iteratively computed considering equilibrium with the room temperature environment and cold stage temperature. A simple model is used to estimate the power sinked to the thermal stages, but this contribution proves to be almost negligible. The model requires a good knowledge of the transmission coefficient of the involved filter layers. They where obtained from several sources: Tucker2006, Ade2006, Benford2003, and Koller2006, along with in-house and third party measurements. After convergence is achieved the model can estimate several parameters of the system. The most relevant for our discussion are the overall transmission and bandwidth of the filter stack, the total amount of radiation that reach the detector, the array noise level, and the main contributors to the noise budget. Table 2 presents the main results from this simulation.
Results indicate that the total load (two bands) on our cold detectors is around 4.5 when observing a 300K load. This value is just within the specification of our sorption cooler. When operating under nominal conditions, i.e. looking at the cold sky, the load diminish to 1.6 .
To estimate instrument sensitivity the sky is modeled as a black body source characterized by its brightness temperature. Power emitted by sky is propagated through the system considering the filters self emission. The noise power reaching the arrays is used to calculate the expected noise equivalent temperature (NET) of the system as 0.8 mK and 1.2 mK for the LFA and HFA. The NET is calculated using the formalism described in Appendix E, including the photon shot noise, bunching and the quasiparticle generation-recombination noise inside the aluminum of the KID. Detector two-level system noise is excluded. Of special interest is the identification of main noise contributors. For the HFA the sky noise dominates the noise budget, meanwhile for the LFA sky and filter contributions are on the same level. Main noise offenders are the input window for the case of the HFA (8.5%) and the optical cavity and cold stop (22%) for the LFA. It is important to note that this estimation just considers the photon noise and KID noise contributions. The measured NET will increase due to the 1/f noise, the read-out noise contribution and other noise sources.
| LFA | HFA | |
| Sky conditions | PWV=1mm | PWV=0.5mm |
| el=60deg | el=60deg | |
| Sky temperature | 70K | 200K |
| Power at array | 0.06 W | 1.57 W |
| Power per detector | 4pW | 102pW |
| NET | 0.82 mK | 1.16 mK |
| Sky contribution | 50% | 79% |
| Stack contribution | 50% | 21% |
8 Read-out system
The read-out system, depicted in Fig. 6, is designed to measure the frequency shift of the KID resonances under nominal load. For this a comb of frequency tones is used, where each tone is tuned to hit a specific KID. By comparing the input and output signals the system computes the complex gain coefficient, S21, associated to each detector tone. This way, small variations on the S21 reflect the frequency and dip depth change of the associated resonator due to the incoming photons. The S21 time streams are stored as raw data in the storage PC using the IQ format: with being the nominal tone power. Simultaneously the system stores the house keeping data (HK), including tone frequency and power among other relevant parameters.
Figure 6 presents a detailed description of the hardware implementation of the read-out system. The architecture uses high speed digital to analog converters (DACs) and digitizer boards (ADCs) to generate and read the required frequency tones for each hardware chain. The procedure can be summarized as follow: Two 14 bit, 2.2 Gsamples/sec DACs in IQ configuration are used to generate an analog 2 GHz bandwidth waveform which is upconverted to the 4-6 GHz range. A second set of DACs on the same board generates an additional 2 GHz bandwidth waveform which is upconverted to the 6-8 GHz range. Both signals are added in a power combiner obtaining a 4-8 GHz analog signal containing a maximum of 1080 tones.
This signal is fed to the KIDs line and the transmission coefficient associated to each detector is imprinted on the output waveform. The signal is amplified by a cryogenic low noise amplifier and delivered to a down-conversion stage located outside the cryostat. Here the signal is down-converted to the 0-2 GHz range by a double sideband mixer. The downconversion process causes folding of the signal around the LO frequency, reducing the required sampling rate of the ADC. The main drawback of this technique is the increase on system noise floor. The possibility of tone collision between upper and lower sidebands is controlled by careful placement of the tones. The three local oscillators (LO) involved in the process are tunable, allowing to adjust for differences in the real pass-band of each detector chain.
The time domain signal is digitized using a 10 bit 4.4 Gsample/sec digitizer, and integrated in the FPGA located in the ADC board. The integration process is done with a buffer size of 216 bins synchronized with the periodicity of the DAC generated signal. As a result a perfect sampling of the waveform is obtained. The resulting integrated signal is send to the back-end computer were a real-to-complex fast Fourier Transform (FFT) is applied. Relevant channels of the spectra, where tone information is located, are selected for further analysis and send to the data merger. The obtained spectral resolution of the system is 67 KHz, enough to place a red-out tone close to the KID resonance (typical width of 300-600 KHz) and to retrieve the relevant information. The back-end computer also controls the waveform generation process. A good synchronization of all digital clocks and analog oscillators allows the extraction of the amplitude and phase of the measured tones, and therefore the calculation of the complex gain coefficient associated to each detector. For this a high-stability and low phase-noise signal generator is used to provide a 100 MHz reference signal to all involved hardware devices.
Each back-end computer is able to process four readout chains, i.e. around 4000 tones. To operate such instrument a cluster of 6 back-end computers is used. All computers send relevant data to the main control computer (KIDDO), where data are merged, formatted, and sent to the storage machine. The maximum speed of the read-out system is around 30 samples per second. The main control computer also does the high-level control of the system, providing a user interface via SCPI commands and configuration files.
9 Data processing and calibration
The readout outputs the complex gain coefficient (S21) of a set of tones placed at the KID resonant frequencies. The S21 time streams need to be converted into the fractional frequency shift of the KID resonance, , which is proportional to the received power signal for small deviations around the average value. This is the case during regular observations as the astronomical signal is a small variation above the atmospheric load. This process requires several steps that are described in this section.
9.1 Tone list generation and blind tone correction
The first step consists of finding the resonant frequencies and command the backend to place the tones there. The backend functionality includes a search routine that performs a full frequency sweep over the 4-8 GHz range (see Fig. 7 for an example). From this data the appropriate frequency and tone power for each KID is determined. This information is written in the tone list file and loaded into the read-out. During this process a set of blind-tones, typically 80 per chain, are distributed over the 4-8 GHz bandwidth in frequencies outside the resonances, as described in Golwala2012. Signals associated to blind tones are stable with respect to the sky signal and measure time variations of the complex gain in the readout hardware itself. Typical contaminating signals are system drifts as amplifier gain variations or thermal drift on cables.
Considering the up and down-converter architecture, described in Fig. 6, blind tones are grouped in quadrants for processing. The data quality of the blind tones is assessed and noisy blind tones are discarded. Selected blind tones are normalized, phase referenced and averaged per quadrant. The result is used as correction factor for the measured S21 time ordered data.
Apart from the correction of all system drifts, this blind tone correction also improves the effective noise floor of the readout electronics. In particular, it effectively removes the large phase noise contribution, of around 30 dB, from the three LOs present in the conversion scheme. As a result, the system achieves a typical noise floor value close to -93 dBc/Hz when operating with 1000 tones. As the typical KID signal is in the range of -90 dBc/Hz we estimate the read-out noise contribution to the instrument NET as an additional 15%, as discussed in Appendix D.
9.2 Circle calibration, frequency sweep and linearization
After blind tone correction, the S21 timelines are converted into frequency changes of the KID resonances (). This process requires that KID resonances are well characterized in the complex plane, i.e. amplitude and phase. This characterization is done via a frequency sweep. It consists of placing the tones of the read-out on each resonance, followed by synchronous sweeping the LOs of the up and down converters in the range from -1.5 to 1.5 MHz. As a consequence read-out tones sweep around its nominal value within the same frequency range. An example of such a measurement is presented in Fig. 8. The result is used to create a full characterization of the KID resonances, which is stored by the analysis software and used to calibrate subsequent regular data. As seen in Fig. 8 resonances in the complex plane corresponds to circles and can be characterized by storing the circle center (Zcenter), the circle radius (), and the phase zero point (). This simple three-parameters model is known as ”KID circle”. For reducing observational data, the blind-tone corrected S21 values are mapped onto the calibration circle using:
Where correspond to the phase shift in the KID circle between the zero-phase point and the current measurement. For small signals on a large background the phase shift is proportional to the fractional frequency shift of the KID resonance, , as shown by Calvo2013 and Bisigello2016. Therefore, it can be used as a measure of the detected power. In the general case the conversion from to is obtained by inverting the relation obtained during the frequency sweep. Fig. 9 demonstrate this relation for a specific resonance. In the plot the linear regime is clearly seen together with the responsivity of the detectors at different tone placement or bias.
9.3 Temperature calibration
To create an absolute temperature calibration scale we use a wire scanner. This is a device that moves an orthogonal pair of 0.5 mm thick wires through the instrument focal plane. The first wire moves along the direction, meanwhile the second scan along the direction. The warm wire movement against the cold sky produces a small signal on the KID detectors that is used to establish positions of each detector in the focal plane, and convert the response of each detector from frequency shift to antenna temperature scale.
The wire size of 0.5 mm was chosen to provide a small signal of around 10 K, assuring that detectors are kept in their linear regime during the measurement. Unfortunately, the required wire diameter is similar to the operation wavelength, 350 m at the HFA and 870 m at the LFA. As a result, diffraction and polarization effects need to be understood (Xie_1991) especially at the low frequency regime, where the wavelength is larger than the wire thickness. Polarization effects average out if we take two orthogonal scan directions, as we usually do to also retrieve detector positions. For diffraction effects we calibrate the effective wire peak temperature by using a wire grid rotator that creates a deterministic absolute temperature scale using two well known calibration sources (liquid nitrogen and ambient load). This information was used to retrieve an effective wire diameter () that accounts for diffraction effects, meanwhile assuming that the wire is in thermal equilibrium with the environment temperature (). As expected diffraction effects are just relevant for longer wavelengths as we find mm for the LFA and mm for the HFA.
Fig. 10 shows the typical time response of a high frequency KID detector to the cable movement. An interesting fact on this signal is that the peak response depends on the optics status: maximum response is obtained when the wire is moved across the focal point of the detector. The key point for calibration is that when the beam is aberrated or defocussed the peak response is lower, but the response is wider. Overall, the integral of the trace remains constant, as long as the detector is operating in its linear regime and no extended response of detectors exist (as seen in Yates_2017).
The wire signal is used to obtain the responsivity of each detector and flat field the camera response. This is implemented in the conversion of data to temperature scale using:
Where the integral in the denominator is the integral strength of the wire response, the physical temperature of the wire, the sky brightness temperature obtained from 183 GHz radiometer measurements, and is the effective size of the wire. Once the detector response is calibrated it is possible to use a part of the timeline as a noise trace (figure 10.b). The noise trace is processed to remove correlated signals and the NET on blank sky is calculated as:
where is the sampling frequency of our read-out.
The wire scanner is used routinely during observations to monitor the status of the instrument and decide the need of a new circle calibration. It is also used to get the optimal read-out power for each resonance: several circle calibrations followed by wire scanner measurements are performed at different read-out power levels. Data is then processed and the individual power setting are chosen to maximize detector sensitivities. We observed that adjacent (in frequency space) KIDs could have differences in power requirements as large as 8 dB, optimal performance is obtained with power adjustment resolution of 0.5dB.
10 Instrument performance
After extensive laboratory characterization, described in Appendix B, AMKID was shipped to Chile and installed in its current configuration in August 2023. The process included the in-situ assembly of the last generation HFA detectors and the room temperature magnetic shield described in AppendixC. After installation several calibration sources were observed using the dual color configuration. These initial observations prove that the observing strategies, pointing model, and beam quality are adequate for both arrays. During the observing season 2025 work focussed on the complete science verification and optimization of the LFA. The HFA verification is still ongoing and final results are expected for end of the 2026 observing campaign.
In the last sections we present the main results obtained during the successful AMKID-LFA commissioning. The current section focus on the measured instrument performance, meanwhile Section 11 presents some early on-sky observations proving the basic instrument capabilities.
10.1 Array sensitivity
The array NET is regularly measured using the wire scanner technique. Fig. 11 shows results of a typical measurement taken during LFA observations. It can be observed that most of the array has an uniform performance around 2.2 (median value). Some detector chip corners show a degraded performance, specially on chip LFA-2. This effect is caused by problems with the gluing process of the lens array. On LFA-1 a small broken corner on the array is seen as an abrupt change in NET value, contrary to the observed degradation in other corners.
For this large number of pixels statistic analysis is mandatory, see Fig. 15. The mode of the distribution represent the performance of well behaving detectors, and can be directly compared with our modeled sensitivity. On other side the median value includes the existence of non-ideal detectors in the array and is therefore a useful value to estimate final performance in terms of mapping speed.
When comparing the measured values of the NET (mode value is 1.7mK) with our predicted values (0.8mK) we observe a large discrepancy. This is further discussed in section 10.2. The main conclusion is that the difference is explained by several factors: the 1/f noise contribution; the existence of contaminating signals; and the digital read-out contribution.
The spread on the NET distribution causes a degradation of the camera performance. In an ideal instrument the median value should approach the mode value as most of the detectors achieve the same performance. The observed spread is caused by several factors such as: the chip performance degradation towards edges, the imperfect biasing of the resonances, and the increased read-out noise contribution at the high frequency range of the read-out. A new generation of read-out technology, currently under development, will help to control the last two factors.
It is important to remark that when initially deploying the instrument the spread of sensitivity was a major problem. The implementation of accurate algorithm for optimizing bias of resonances caused a reduction of 20% in the NET median value.
10.2 Stability
System stability is studied by analyzing the power spectra of a 300 s noise scan done on blank sky with a static telescope. The noise trace is then decorrelated to remove atmospheric fluctuations and analyzed. The Fig. 12 presents the noise power spectrum of four representative detectors, one per sub-array, with a NET near the peak of the distribution. To facilitate visualization the power spectra is smooth to 0.1 Hz. The figure shows the traces before and after atmospheric noise removal. Results indicates a 1/f knee point of 0.5 Hz, identical to the value reported by the independent pre-integration chip level measurement discussed in section 4.
The full resolution spectra allows us to disentangle between the different noise sources affecting detector sensitivity. Due to the 1/f contribution the measured NET is larger than the noise level of the plateau above 2 Hz. This last value, indicated with a green line, is expected to be comparable with the photon-noise level predicted by our simulations. This 1/f contribution arises from two sources, the inherent detector 1/f noise plus the residuals from sky fluctuations removal.
A second relevant source of noise is the existence of unwanted contaminating signals, mainly at 0.3 Hz and harmonics up to the maximum frequency. To estimate its relevance we remove them from the spectra by using a median filter on the power spectral density before the smoothing process. The result, shown as a red line in Fig. 12 represents the estimation of the photon noise plus detector noise for each sub-array.
The above analysis indicates that we have a contribution of around 10% to the NET budget from contaminating signals in the spectra and 15% contribution from the 1/f noise. Additionally we have to account for a 15% degradation due to the read-out contribution. Values discussed here have a large spread as many factors add up. As an example, detectors that have a readout frequency in the high band have a larger noise contribution from digitization process than low band detectors. Or the fact that the 0.3 Hz signal is not uniform over the array, being much stronger towards the edges of the detector chips. With these considerations we conclude that the photon noise level of our detectors accounts for almost half of the measured noise.
This exercise is fundamental for instrument development, as it shows the path for further improvements during the coming years. Of special relevance is the deployment of a new generation read-out that holds the promise of lowering the digitization noise contribution and diminish the observed NET spread. Another possible improvement is the suppression of the 0.3 Hz noise. This noise signal was initially observed during laboratory testing of the instrument, see Appendix C, as a strong feature in the timelines. The addition of a magnetic shielding to the instrument attenuated the signal around 40 times, but as discussed above it is still present on the system. Although the exact source of the signal is still unknown, the signal could be further attenuated by improving the magnetic shield of the instrument.
10.3 Optics performance
Early results show that the instrument optic performs according to expectations. Planet scans, as the one presented in Fig. 13 are used regularly to measure beam properties. Average beam sizes of 17.5” with beam ellipticity lower than 1.1 are regularly measured with the LFA.
A relevant optical parameter is the detector coupling to a point source (), i.e. the conversion factor from temperature scale to source flux in Jansky. It is defined as:
Where is the telescope collecting area, kb the Boltzman constant, and the aperture efficiency of the instrument and telescope, including Ruze efficiency. The factor two comes from the fact of total flux being in double polarization mode. Optical coupling is simply measured by comparing the expected flux from a calibrator, usually a small planet as Mars in opposition or Uranus, to the measured KID response in temperature scale. As expressed in Eq. (4) it corresponds to a direct measurement of the instrument aperture efficiency. Measurements, presented in Fig. 14 and 15, indicate an average optical coupling of Jy/K at 870 corresponding to an instrument aperture efficiency of 61%. It is expected that the current efforts to improve the optical quality of the system, characterized by a half wavefront error of 26 m RMS (see Reyes_2023), will further improve this value to around Jy/K.
In Figure 14 the distribution of optical coupling along the field of view is presented in color code. Measured beam sizes are shown as ellipses sizes. The LFA presents a uniform response over the array with some edge degradation, especially on chip LFA-2 which are currently under study. The performance was initially measured after instrument installation and proves to be stable over already a calendar year.
10.4 Detector yield
The first step for instrument operation is to perform a KID search and identify the resonances that can be used for observations. The list of resonances is stored by the back-end and used in subsequent observations. The first definition of detector yield is therefore how many of the existing detectors were correctly identified. Main limitation is the fact of resonances not being equispaced in frequency and tending to cluster as shown in Fig. 7. When two resonances are located at a distance lower than a resonance width they start to interact and exhibit a coupled response as shown in Bisigello2016SPIE. In extreme cases resonances are so close together that they cannot be distinguished given our read-out frequency resolution. Our algorithm uses a threshold of 0.15 MHz placing a tone in only one of the two resonances when a collision is detected. Also cross-talking pixels that survive this first filter will give deteriorated NET and beamshape, so will be filtered out during next steps.
After correct identification of resonances the wire scanner is used to get an initial estimate of beam positions on sky and NET. Detectors with sensitivities larger than a certain threshold are not useful and will be removed. This causes a second reduction in yield. Note that, as wire measurement is insensitive to beam shape, we reject here pixels that have a poor performance mostly caused at chip level. Main offender is the degraded performance of lens array towards the edge of some sub-arrays, the existence of crosstalking detectors not removed in previous step, and the read-out contribution that increases towards high readout frequency range.
While the wire scanner position provide a first estimate of the detector position within the array, their information is not precise enough for astronomical observations. Accurate positions are determined from on-the-fly (OTF) maps across Mars or Uranus with parameters set such that we obtain a fully sample map of
the planet for each detector (beam maps). For these maps not only the information about the relative position for each detector is obtained, but also the optical coupling of each detector is measured by comparing the expected planet flux to the measured temperature calibrated by the wire scanner and the beam ellipticity.
The Noise Equivalent Flux Density (NEFD) of each detector can be estimated by the simple multiplication of the NET and measured optical coupling. Results allow to derive the final yield of the instrument as detectors that have a sensitivity in the target range. Results, presented in Table 3 show that the final on-sky yield of the LFA is around 71%.
| Array | LFA |
| Number of pixels | 3520 |
| Detected resonances | 3030 (86%) |
| NET in range | 2568 (73.0%) |
| NEFD in range | 2487 (71.0%) |
11 Instrument early results
Several verification and early science observations have been carried out during the 2024 and 2025 observing seasons. The campaign verified that under good atmospheric conditions the calibration routine is stable and allows to observe for several hours without repeating the frequency sweep calibration procedure. As good atmospheric conditions we consider a PWV lower than 1.0 mm for the LFA. The need for recalibration is asset by regular wire scanners and by regular observation of calibration sources to monitor the instrument sensitivity. Our observing strategies rely heavily on those developed for LABOCA (Siringo2009) in the past. The strategies include the mapping routines to determine the telescope foci, as well as the spiral and raster scanning patterns. We corroborate that determining the focus on the central area of the array is representative and no degradation is observed due to focus settings towards the edges of the camera. A telescope pointing campaign using the HFA, demonstrates that the pointing model is stable, and no major correction is required due to the tilting in elevation of our massive optic system.
After these initial verifications done with both arrays we focused our efforts on the commissioning of the LFA. That array is currently validated and ready for science operation. Flux calibration based on the wire scanner data to obtain the gain of individual pixels plus the determination of optical coupling with planet scans proves to be successful obtaining a typical calibration error below 10%. Sensitivity verification using calibration sources show that on sky sensitivities are in the range of 70-90 mJy, matching the predictions from NET and optical coupling measurements.
The final RMS noise () when mapping a square area of size and with an integration time of can be calculated using:
where is the median of the measured NEFD, is the effective pixel number of the instrument, and Tatm is the atmospheric transmission. The array size accounts for the need of extending the area to be mapped to get homogenous noise coverage at the edges of the map. The sampling factor () is usually set to two to obtain Nyquist sampled maps. Finally the beam area () can be easily calculated from instrument nominal beam size (17”). Using this formula the instrument mapping speed can be calculated as 25mJy (including 25% overhead) per square degree in an hour integration when operating under nominal weather conditions of 0.5mm PWV.
As an example we show in Fig 16 an image of the nearby (3.8 Mpc) starburst galaxy NGC4945 obtained in 2 hours on source integration time (compare with Weiss_2008)) as well as the high-redshift dusty starforming galaxy SPT0311-58 at z=6.9 (Strandet_2017) with an on-source integration time of 2 hours. Results indicate that the AMKID-LFA is 3.8 times faster as LABOCA at APEX when mapping extended sources. When comparing both instruments it is important to consider the different instrument architectures, as while AMKID is a single polarization oversampled array (1.2F) covering a square field-of-view of 15’, LABOCA was a dual polarization horn coupled array (2.5F) with 290 detectors over a circular 12’ field-of-view. For discussion on the subject the reader can refer to Griffin_2002.
The AMKID data can be processed using the MARS software package. MARS builds up on the Bolometer array data Analysis (BOA) software package developed for LABOCA data reduction, but its design has been changed to support parallel data processing on multi-core machines as well as including all specific AMKID reduction and calibration steps described here.
12 Conclusions
The revised AMKID camera was successfully installed at the APEX telescope. Performing with an unprecedented number of pixels, dual color observing capabilities, high sensitivity, large field-of-view, and high angular resolution, this instrument is expected to open a completely new range of science with APEX in the upcoming years. In this paper we present an overview of the instrument design with a special focus on the science commissioning of its low frequency channel. Early results indicate that the camera achieves a typical sensitivity per beam at 350 GHz around 70-90 mJy per beam.
The dual color operation of AMKID offers an additional high frequency channel at 850 GHz. Initial observations using the AMKID-HFA demonstrate the high frequency channel capabilities, but also highlighted the need for higher quality optical components to start successful scientific operation. The new optic system, consisting of a new set of warm mirrors with a lower RMS error than current system, will be installed soon. With this improvement HFA early science results are expected by the end of the 2026 submillimeter observing season. On other side a new read-out, currently under development, hold the promise of an overall improvement in sensitivity thanks to the lower read-out noise contribution but also due to finer tone placement, therefore improving current detector yield.
Acknowledgements.
The Authors thank T.M. Klapwijk for his work on the early development of the Aluminium-NbTin hybrid KIDs.This work is dedicated to the memory of Karl M. Menten (1957–2024) whose enthusiastic support of the APEX project made the development of AMKID possible.
Appendix A KID detectors fabrication
The AMKID detectors are antenna-coupled coplanar waveguide (CPW) resonators made of a 400 nm film of NbTiN with a K as described in Jansen_2013. The exact geometry is shown in Figure 2 with dimension specified in Table 4. They are fabricated on a high resistivity silicon wafer, as described in Thoen_2017 and Bos_2017. This section give some further details on our detectors design and fabrication.
At the open end the resonator is coupled to the readout line, here the CPW line is wide to reduce 1/f noise Gao2008. At the resonator shorted end there is a section where the CPW line is much narrower with the central line made from aluminum. The length of this section is 30% of the KID length, i.e. about 1-2 mm long depending on the KID resonant frequency. This scaling improves the readout frequency scatter of the KIDs, since any variations in aluminum line width will have the same fractional effect on all detectors. Another advantage is the frequency independent responsivity and photon noise level.
The resonator is coupled to a twin-slot antenna (see Zmuidzinas_1992; Filipovic1993), which couples the radiation from the telescope into the NbTiN-Aluminium hybrid CPW by means of a Silicon lens mounted on the chip backside (see Fig.2b). Quasiparticle creation only occurs in the aluminum line, since the gap frequency of NbTiN is 1.1 THz and thus provides a completely loss-free material for the resonator and antenna structures.
Both high and low frequency detectors are very similar, being the main difference the antenna dimensions which scale with operating frequency.
| Antenna | h | a | b | w |
| LFA | 240 | 12 | 12 | 25 |
| HFA | 97.7 | 6.0 | 2 | 55.8 |
| Lens | ||||
| LFA | 0.9901 mm | 0.3861 mm | 1.90 mm | 2.00 mm |
| HFA | 0.5517 mm | 0.2152 mm | 0.940 mm | 1.00 mm |
| KID | |||||
| LFA | 1.6 | 2.2 | 14 | 24 | 40 nm |
| HFA | 1.6 | 2.2 | 6 | 10 | 65 nm |
AMKID operates under very significant loading powers, of about 4 pW for LFA and 100 pW for HFA as discussed in Table 2. As a result the KID internal quality factor (), is strongly suppressed, even more for high frequency resonators with a smaller aluminum volume. The KID dip depth () is related to the KID coupling factor () and as:
Since it is crucial that the KID dip depth is between -3dB (0.5) and -6dB(0.75) to ensure a low enough noise contribution from the electronics, and is found to scale with KID frequency we scale similarly.
Additionally, we design the KID resonant frequencies such that the KID to KID frequency spacing increases linearly with frequency, i.e. is constant, to maximize the packing of the resonators. The frequency encoding on the chip is that nearest frequency neighbors are always separated by one detector at a completely different frequency Yates2014. This prevents electromagnetic cross-coupling and improves the frequency scatter, as spatial variations of material parameters vary slowly over medium to large distances.
Appendix B Laboratory instrument characterization
Before deployment at the telescope, the instrument was extensively tested in laboratory environment. Performed measurements and analysis include optical quality and alignment, instrument sensitivity, and stability, among others. A dedicated experimental setup was built to mimic the instrument operational conditions. The system includes a mechanical structure to tilt the instrument, equivalent to 60-90 deg of elevation at telescope; a wire scanner to generate small signals over a cold background load;and a broad band photonic continuous wave signal generator mounted on a X-Y stage allowing for beam characterization of the system, as described in Reyes_2023.
Of particular interest is the wire scanner device, presented in section 9. It consist of two small wires that moves against the sky background. Position of the wires is accurately record using Precision Time Protocol (PTP) over Ethernet and synchronized with data acquisition. The wire size (0.5mm) and material(standard electronic wire, plastic coated) were chosen to generate a small 10 K signal against the cold background.
The wire scanner was originally developed for fast characterization of the instrument performance in laboratory enviroment. For this a cold load is placed behind the wires mimicking the sky background. The load wass implemented as a large copper bath covering the complete focal plane (80x80 cm2). The bucket lower side is externally coated with Stycast and silicon carbide grains to act as a cold load. The bath is thermally insulated by placing it inside a larger 4 cm wall thickness styrofoam box. The effective temperature of this cold load was measured using a KID detector together with a rotating grid. The grid switches between two standard temperatures, ambient load at 300 K and liquid nitrogen load at 80 K. Measurements indicate an effective temperature of the load of 120 K for the low frequency band 160 K for the high frequency band, elevated from 77K due to the styrofoam losses. However, these values are similar to the expected sky temperature at the deployment site.
The wire-scanner characterization method proves to be effective and very fast. It’s main purpose is to measure the temperature scale conversion factor, and give a first estimation of detector sensitivities. Even when not perfect, several optical parameters as beam sizes and ellipticity, can be also estimated from the wire response shape. The complete array can be characterized in a couple of minutes; as a consequence, this method was extensively used as a basic characterization routine during the instrument optimization phase. Several optical issues were detected and corrected using this method.
Appendix C Magnetic shielding
During instrument optimization it was clear that AMKID was extremely sensitive to external magnetic fields. This was specially prominent on the LFA, possibly due to the larger pixel size, almost 10 times larger than HFA pixels, enhancing the coupling of the detector to environmental magnetic flux. This phenomena was seen on various aspects of instrument operations:
Trapped flux during cool-down makes the resonances frequency to significantly move from one cool-down iteration to the next one.
The low frequency array presented a prominent 0.3 Hz signal over most of the array, but specially strong at chip edges. The amplitud of the signal was as large as 0.5 K. This signal is most likely magnetically coupled to the detectors.
Initial testing of the instrument at the telescope in 2016 showed a strong signal of rad (around 5 K) correlated with telescope azimuth movements. This was caused by the varying coupling to the earth magnetic field.
Visible KID response is obtained by waving a ferromagnetic material (scissor, screwdrivers) around the cryostat.
This behavior presented a challenge for operations, but as induced signals are slow and correlated over the array they can be partially filtered out in post-processing. A more serious problem is the hysteretic response of some detectors. This is seen as a sudden jump of level of the timelines of some detectors, usually happening when magnetic flux exceeds a threshold as described in Stan_2004. This high frequency content signal cannot be filtered out and the affected detectors cannot be used for astronomical observations. This topic was investigated experimentally in laboratory environment by mounting an external small coil on the cryostat ( see Fig. 17) to artificially generate magnetic field as expected in regular operations. The main conclusion of the study is that the HFA performs well when facing magnetic fields up to 15 T (maximal change during an Azimuth scan due to earth magnetic field). For the same field intensity the LFA presents a hysteric behavior. This noticeable difference is due to the increased pixel size at low frequency, giving a larger focusing of flux into the sensitive part of the KIDs. Other possible reasons for this difference are the higher kinetic inductance of the LFA KIDs due to the thinner Aluminum film, and the slightly worst shielding of the LFA by the Aluminum cup.
The AMKID original design included some magnetic shield consisting of an Aluminum superconductive cup, with a moderate aspect ratio of four, located around each of the two arrays. Additionally, the detector chips include a superconductive back-plane which also act as a shield for detectors located at center of chip, but deteriorating edge detectors. The laboratory experiment shows that this shield was not sufficient and that a factor 40x of additional magnetic shielding was required to avoid jumps in timelines. This was addressed by including a passive magnetic shield around the cryostat outside. It consists of a series of 1 mm thickness plates made of high magnetic permittivity metal ( around 50000), covering a large fraction of the cryostat. The coverage was carefully designed using numeric simulations for easiness of installation, meanwhile achieving the required attenuation at all possible orientations of the cryostat. After shield installation the above mentioned affects are no longer visible and a typical external field attenuation of 100 is observed during operations.
Appendix D The readout noise contribution
The digitization noise caused by the finite number of bits in the ADC and DAC can be expressed by the power spectral density (PSD) of the read tone normalized by the off resonance tone power as:
where GSamples/s is the sampling rate of the ADC; is the noise degradation due to the number of tones (n) used by the read-out; is the crest factor, the ratio between the peak voltage and the root mean square voltage of the waveform. We use a random initialization of tone phases achieving a typical crest factor of 14 dB. The factor 3 dB is because of the the double sideband down converter, which increases the noise floor by a factor of 2. The effective number of bits (ENOB) is a measure of the real performance of the digitization hardware and is lower than the real number of bits. The AMKID read-out uses a 14 bit DAC and a 10 bit ADC, being the total PSD of the system dominated by the ADC noise, characterized by an ENOB of 7.6dB.
For 1000 tones the read-out achieves a of -93.8 dBc/Hz, assuming flat power across the band.
The system noise analysis shows that our read-out noise is dominated by the above described digitization noise and not by the analog section of the read-out, being the last more than 2 order of magnitude smaller that the digitization noise (-125.2 dBc/Hz).
The digitized signal is integrated in a circular buffer and later Fourier transform as explained in section 8. The output signal from the digital section is the gain coefficient associate to each readout tone. We can express the digitization noise as a standard deviation of the measured gain coefficient in a one second measurement as.
Where is the channel width, the number of integrations in the circular buffer, and the number of samples per second. The above expression can be further simplified as . The measured standard deviation of the S21 can be temperature calibrated following the procedure described in Sec. 9 and then expressed as a contribution to the system NET:
Where K is the conversion factor from phase shift () to temperature scale as discussed in section 9; fsamp the sampling frequency; and R is the radius of the KID circle, which is related to the KID depth (min(S) as:
During observations the noise performance of the system is regularly monitor by executing noise scans after a calibration sweep. This data is analyzed and the measured as the standard deviation of the data along the KID circle radial direction as shown in Fig. 19. At the same time the standard deviation along the tangent direction is computed as . The measured NET is proportional to , meanwhile the digitization NET contribution is directly proportional to the value. This fact is used to estimate the relevance of the NET contribution to the total measured NET using the following expressions:
In our system we obtain a typical separation (ratio between and ) of 2.8 dB, meaning that the ADC contribution is on the same order of magnitude than the instrument NET. Giving the cuadratic adding of NET this means that the ADC noise accounts for a 15% of the measured NET, indicating a clear path to improve the system performance by using a better read-out with a larger ENOB.
Appendix E NET modelling
The amount of power per unit of frequency that reach the detector, , can be divided in two main sources: The power arriving from the sky background, , and the power that is generated by the filter stack itself, . As filter stack we consider all elements that are placed from the cryostat window to the last bandpass filter, including the polarization grid and the cold-stop. Note that the lens array is not considered as part of the filter stack.
The term is calculated using the contribution from each element in the filter stack. For this an iterative model is used: all elements are modeled by its physical temperature and its transmission, reflection and absorption coefficients. At each iteration the emitted, absorbed, reflected and transmitted power are computed for each filter and used to calculate the thermal balance of the system when exposed to the ambient temperature. The process is repeated until the system reach thermal equilibrium. As a result we obtain the physical temperature and emitted power of each filter. This information is used to finally obtain .
On the other side, corresponds to a black body spectra with an associated temperature which is filtered by the filter stack. It is important to notice that both terms, and , include in one way or another the polarization efficiency (=0.5), the cold stop efficiency (), and the window and filter physical sizes. They also include the transmission of all other elements in the stack, which for convenience we group in a term called .
To obtain the power that reaches an individual detector we use the following expression:
Which calculates the power that reach each detector as the ratio between the detector area () to the full array area ( m2) multiplied by the lens-antenna radiation efficiency () and the coupling of the lens-antenna to a plane wave (86%). This last factor is used as our lens-antenna system produces a gaussian electric field distribution at lens aperture and not an uniform distribution as expected for noise photons.
The number of photons per unit of frequency that reach a detector during a second is calculated as:
From here we calculate the Noise Equivalent power (NEP) of a single detector as :
Expression that include the Poisson noise together with the wave and recombination noise contributions, see Flanigan_2016 and Ferrari_2018a. is the efficiency of the quasiparticle generation and is the superconductor bandgap frequency. The factor is used to express the NEP in units of , i.e. as the S/N obtained in a 0.5 s integration time.
As last stage we convert the NEP measured at detector level to NET measured at instrument focal plane, value that is comparable with our measured NET using the wire scanner.
Where is the Boltzmann constant and the filter bandpass. The factor is used here to express the NET in units of , i.e. as S/N after a 1 second integration time. In our case the factor 2 in the denominator cancels with the , but it is explicit in the equation to facilitate extension of this procedure for dual polarization instruments.
All the values used in the above calculations had been discussed in the article main body. To facilitate the use of above equations we provide table 5 showing a summary of the required values.
| Parameter | LFA | HFA |
| NET | 0.83 mK | 1.16 mK |
| 1.0 mm | 0.55 mm | |
| 0.84 | 0.76 | |
| 0.45 | 0.55 | |
| 0.5 | 0.5 | |
| 0.56 | 0.41 | |
| 345GHz | 850GHz | |
| 38 GHz | 123 GHz |